If your front lawn is 23.0 feet wide and 20.0 feet long, and each square foot of lawn accumulates 1050 new snow flakes every minute, how much snow (in kilograms) accumulates on your lawn per hour? Assume an average snow flake has a mass of 1.80 mg. Number kg/ h
Answer #1

Solution)

Here,

Area = 23.0 feet * 20.0 feet

= 460 feet^2

Now,

number of snow flakes per minute = Area & number of flakes per unit area

= 460 * (1050)

= 483000 snowflakes

Now, number of snow flakes in 1 hour = 483000*60 snowflakes

= 28980000

Here, mass of 1 snow flake = 1.60 mg = 1.60*10^-6 Kg (Given)

So,

total mass = 28980000 snowflakes * 1.60*10^-6 Kg/snowflakes

= 46.368 Kg (Ans)

==========

Good luck!:)